# Longitudinal Abuse and Sentiment Analysis of Hollywood Movie Dialogues using LLMs

### v2

@@ 3.2 N-gram analysis
The N-gram model is a basic NLP method used to examine ”n” consecutive words from a given sequence of text [65]. The N-gram model has been widely used in text categorisation and provides further insights about sentiment classification [66] and useful and improves the classification performance of informal text-based tasks [67]. The simplicity and efficiency of the N-gram model make it the preferred choice for large vocabulary speech recognition applications [68]. Pang et al. [69] employed unigram, bigram, and a combination of both unigram and bigram techniques to classify positive and negative sentiments for classification. In our framework, we utilise bigrams and trigrams to ascertain the frequency of word occurrences, combined with sentiment analysis.
3.3 Deep learning-based language models
The LSTM network has had a significant contribution to NLP tasks including machine translation, speech-to-text transcription, and language modelling applications [70, 71]. NLP strategies typically employ word embedding models to transform text into numeric representation for deep learning models. Word to Vector(Word2Vec), Global Vector(GloVe), Embeddings from Language Models (ELMo), and BERT are commonly used word-embedding [72][73]. ELMo [74] is a context-based word embedding method. that learns contextualised word representations based on a neurolinguistic model based on a character encoding layer and two Bidirectional-LSTM layers. BERT is a pre-trained language model utilising bidirectional representation for masked word embedding based on context. In contrast to GPT models, BERT incorporates bidirectional contextual and is often used for word embedding or refined as a model using task-specific data. The original BERT consists of two strategies for applying pre-trained models to downstream tasks: feature-based with 12 encoders and fine-tuning with 24 encoders [75], ELMo belonged to the former, and GPT was a type of fine-tuning. The RoBERTa (Robust BERT) [76] is an optimised BERT model utilising a larger data corpus for pre-training. It can process longer text sequences, and dynamically change the masking pattern of the training data [77]. RoBERTa uses more parameters and a deeper network structure during pre-training. This enhancement significantly improves its performance in downstream tasks. The pre-training phase enables the BERT and RoBERTa models to acquire substantial knowledge, thereby facilitating its ability to effectively handle the target task with limited samples during the fine-tuning phase [78]. In our study, which comprises movie abuse and sentiment analysis, fine-tuning such models with appropriate data is critical.
HateBERT [34], a variant of BERT with a deeper architecture and more parameters, enhances linguistic representation learning, similar to RoBERTa. This design improves its performance in hate speech detection tasks. Pre-trained with robust language comprehension capabilities, HateBERT adapts quickly to diverse downstream tasks during fine-tuning, even with limited labelled data. Our research shows that fine-tuning HateBERT on domain-specific data enables effective detection of abusive language in films and improves fine-grained sentiment analysis. This domain-specific training helps the model capture unique linguistic patterns, boosting its practical accuracy.
3.3.1 SenWave dataset
The SenWave [42] dataset for sentiment analysis comprises over 105 million collected tweets worldwide, providing an assessment of global sentiment fluctuations during the COVID-19 pandemic. This comprehensive dataset includes annotations for 10,000 English and 10,000 Arabic tweets across ten distinct categories: optimism, gratitude, compassion, pessimism, anxiety, sadness, anger, denial, official reports, and jokes. Qiang et al. [79] preprocessed the raw data by filtering irrelevant characters using the NLTK (Natural Language Toolkit) tool. The SenWave dataset has been used for identifying trends in the change of sentiment for post-COVID-19 studies [75] and also used for refined pre-trained language models for machine translation analysis for low resource languages [80]. These studies motivate the usage of the SenWave data for analysis of Hollywood movie dialogues.
3.3.2 RAL-E dataset
RAL-E (Reddit Abusive Language English) [34] is a large-scale dataset collected 1,478,348 messages from the Reddit 777www.reddit.com online communities called Subreddits [81] that were banned for offensive, hateful, and abusive content. The Subreddits were created and controlled by users, so moderators could dictate the rules for their Subreddits [82]. In June 2015, Reddit banned hate speech-related Subreddits for violating the site’s user agreement and tightened its content policy [83, 34]. Tommaso et al. [34] provided an analysis of the abusive vocabulary of the Subreddits and compiled a dataset enabling further research. The RAL-E dataset has been utilised to train several transformer-based deep learning models, which have shown impressive performance in detecting online abuse [34, 84, 85]. In this study, we use the RAL-E dataset to fine-tune the HateBERT model to improve its abuse detection capabilities.
@@ 3.4 Framework
We present a framework for abuse detection and sentiment analysis of movie subtitles that includes several sequential steps, as shown in Figure 1. Our framework uses three datasets, including the movie subtitles, the SenWave dataset for sentiment analysis and RAL-E dataset for abuse detection for refining the RoBERTa models.
In Stage 1, we download and curate movie subtitles, where we considered the Oscar-nominated films and the top 10 films at the box office to be the most representative films of each year with categories shown in Table 1. In Stage 2, we use NLTK [79] tool to remove the special phrases, such as hashtags, emotion symbols (emojis), stop words (e.g. “the”, “an”, “you”) for each movie subtitles during data processing. We need accurate bigrams and trigrams for subsequent analysis, so we also eliminated stop words based on specific categories. In Stage 3, we analyse the processed movie subtitles and extract the top ten bigrams and trigrams for longitudinal analysis. We also report movie duration changes over the decades. We also process and clean  the RAL-E dataset and the SenWave dataset for fine-tuning the BERT-based model at a later stage.
In Stage 4, we use two RoBERTa models obtained from HuggingFace [86] 888 https://huggingface.co/docs/transformers/en/model_doc/roberta and fine-tune them with SenWave and RAL-E datasets for sentiment analysis and abuse detection, respectively. The SenWave dataset enables comprehensive sentiment analysis by providing multi-label classification.
In Stage 5, we categorise the sentiments based on abuse analysis and sentiment analysis. With the abuse analysis, abuse emotions were categorised such as profanity, porn, insult and spam. Our sentiment analysis provides classifications such as denial, pessimistic, thankful, and optimistic to analyze positive and negative emotions more precisely. In Stage 6, we analyze movie subtitles from different eras and categories using data visualisation taking into account Oscar and Blockbuster categories along with the specific groups (action, comedy, drama, horror). We compare the percentage of abusive text in each category and longitudinal sentiment polarity. and then more comprehensively compared a wider range of different emotions, both positive and negative. Finally, in Stage 7, we select popular Oscar and Blockbuster movies, provide sentiment analysis and abuse detection on the entire movie timeline, and review it by taking snapshots of dialogues for qualitative analysis.
@@ 3.5 Technical Details
We utilise state-of-the-art NLP techniques with pre-trained models based on RoBERTa  [^9] https://huggingface.co/FacebookAI/roberta-base, specifically the hateBERT model [34] optimised for detecting hate speech. We used the pre-trained HateBERT models via the Transformers library  [^10] https://github.com/huggingface/transformers. Given the large dataset, we implemented batch processing to efficiently make predictions.
Initially, we implemented BERT and RoBERTa models and found that RoBERTa was more efficient in refining using the SenWave dataset.
@@ 4.1 Data analysis
We use N-gram analysis to outline the changes in the top twenty words of theme and sentiment in selected films over every 20 years. In Figure 2-Panel (a), we can observe that from 1950 to 1969, positive and colloquial phrases such as ”yes sir”, ”let go” and ”good night” were frequent, and these words usually appear in contexts that express affirmation or goodwill. In addition, ”ha ha ha” and “good night good” indicate that the movies from this period are positive and affirmative emotions. Between 1970 and 1989 (Figure 2)-Panel (b), we can see that common bigrams such as ”let go” and ”yes sir” still have a dominant position and ”oh god” and ”come come” have risen significantly. These changes reflect shock, surprise and action in the film. The appearance of ”come let go” and ”go go go” shows an increase in the film’s depiction of the action. In the period 1990-2009 (Figure 2-Panel c), we can see that ”let go”, ”oh god” and ”come come” still appear more frequently in bigrams than previously. The common trigrams such as ”go go go” and ”come let go” continue to appear, but ”hey hey hey” rises in frequency. The rapid increase in word frequency indicates a significant increase in the length of subtitles in movies over the past two decades. The similarity of high-frequency words indicate that the main emotions of films during this period are the same as those of different periods in the past. During the period 2010-2024 (Figure 2-Panel d),, both bigrams and trigrams contain high-frequency particles, such as ”hey hey”, ”yeah yeah”, ”whoa whoa whoa” and ”oh oh oh”. This indicates that the emotions in films during this period were still predominantly positive, accompanied by more vocabulary that enhanced emotional expression.
@@ 4.2 Sentiment analysis
Movies are generally highly expressive, and there is often no single emotion that can be expressed in a film. The different types of emotions captured from the subtitles depend on the type of film as well as the specific plot. Figure 4 illustrates the frequency of different emotions in movies detected by RoBERTa refined using SenWave data. We find that humour (joking) is the most frequent sentiment, which indicates that humour and lighthearted elements are essential in engaging audiences. This is followed by ”annoyed”, ”sad”, ”pessimistic”, and optimistic sentiments which include all the selected movies from 1950 to 2024. We note that these include Oscar-nominated movies and blockbusters in four major categories. Figure 5 presents the results for the movies in the four major including action, comedy, drama and thriller. We can observe that the humour sentiment is expressed the most in all categories, followed by ”annoyed”. The proposition of ”sad” is different, where drama has the most and action and thriller have the least. We can observe that there is a high number of ”sad” sentiments in the ”comedy” category. Most movies would fall into two categories. For example, the movie ”Annie Hall” is an example of a movie that would fall both in comedy and drama and ”The Departed” would fall both in action and thriller categories. Therefore, the sentiments detected have overlaps across the categories.
Figure 4: Counts of different emotions for Oscar nominated and top ten blockbuster movies from 1950-2024.
@@
We input the entire movie subtitle dataset into the RoBERTa model to generate word embeddings. A fully connected layer was then appended to the RoBERTa output, with the output dimensionality corresponding to the number of sentiment dimensions. Using an unconstrained linear activation function, we acquired sentiment weights ranging from -4 to 3, as depicted in Figure 6. This demonstrates the contribution of different emotions to the overall affective polarity score, where positive numbers indicate positive emotions and negative numbers indicate negative emotions.
Based on Figure 6, we calculated the Average Sentiment Weight score for every decade as shown in Figure 7. We observe that for blockbuster movies, the weighted sentiment scores are much higher during the 1950s, 1980s, and 1990s when compared to other periods, indicating that movies had richer emotional expressions during these three decades. In the case of Oscar-nominated movies, the average weighted sentiment score has been much higher than other periods since 2020, indicating that the Oscars are gradually increasing their focus on films with richer emotional expressions. The standard deviation values in the bar chart are substantially higher during the 1950s, 1990s, and 2000s compared to other periods, indicating notable differences in emotional expression among films in these three periods.
@@
Figure 7: Average Weight Sentiment Score Trend Over Years
Different themes and styles of movies cover different emotions, and multiple emotions can coexist in the same movie.Therefore, we obtain a heatmap of the co-occurrence of emotions in Figure 3, where the darker colours show higher co-occurrence frequency. We can observe that ’Humour’ and ’Annoyed’ have a high frequency that suggests many films use humour alongside anger, e.g. situations where jokes cause discomfort. Often, satire and dark humour are employed to show societal harsh realities. Similarly, ’Joking’ and ’Sad’ co-occurred 97 times, indicating that humour is often used to alleviate sadness, a characteristic commonly observed in tragicomic films such as the infa.
At the same time, there are some low-frequency co-occurrence emotion phrases that deserve our attention. ”Thankful” and ”Homour” are shared only once, which rarely go hand in hand in movies. The co-occurrence frequency of ”Thankful” and ”Pessimistic” is 0, which indicates that these two emotions hardly appear and in reality, thankful and pessimism are almost two opposite emotions and rarely appear in the same context.
@@ 4.3 Abuse and Hate-Speech detection
We next use HateBERT for abuse and hate speech detection, which is derived from the pre-trained RoBERTa model and fine-tuned using the RAL-E dataset. Figures 8 and 9 present the results for abusive content, demonstrating how the movie subtitles have changed over the years influenced by social-cultural, censorship, film genres, and external factors such as epidemics.
We utilise a look-up dictionary to count the number of abusive words. Figure 8 shows the trend of the number of words and abusive words in the film subtitles for each decade. We observe that the trend of the number of total words and the number of abusive words exhibits a greater degree of similarity across different decades. After the 1970s, the proportion of abusive words to total words in film subtitles across decades has not changed significantly. Such fluctuations may be attributed to changes in social tolerance for film content, the influence of censorship, and the emergence of new film genres.
@@
Figure 9: Average movie time vs average abusive movie time per decade.
@@
Figure 10 presents trends in the amount of violent content detected in Oscar-nominated films and top 10 blockbuster films between 1950 and 2024. We observe that abusive content has shown a gradual increase since 1980, peaking in 2014 which may reflect the growth of the film industry and the changing social and cultural acceptance of abusive content. From 1950 to 1980, the amount of abuse in Oscar-nominated and top 10 blockbuster films was relatively small. This could be due to more conservative social and cultural requirements of film content at the time. However, abusive content gradually increased after the 1980s and the amount of abuse in the top 10 blockbuster films almost surpassed that in Oscar-nominated movies before 2010, and fallen significantly since then. This may indicate that Oscar-nominated films have become bolder and more diverse in subject and content, and the increase in abusive content may be an attempt to better represent complex social issues and human conflicts. Overall, we observe evolving trends in the representation and amount of abusive content in films over time, which reflect changes in social culture, the film industry, and audience preferences. Before 1980, abusive content was less and more restrained, and then increased with the diversification of film genres, especially in commercial films. Abusiveness in Oscar-nominated films has increased significantly since 2010, reflecting films’ exploration of complex themes and commercialisation.
We selected the years with the highest and lowest number of abuse for Oscars and Blockbuster movies in Figure 10, and displayed the corresponding movie titles in Table 2. Analysis of Oscar-nominated films shows a peak in abusive content in 2014 and 2020. Films like The Wolf of Wall Street and The Irishman use frequent expletives to heighten dramatic tension, emphasise the protagonists’ volatile personalities, and create a gritty atmosphere. 12 Years a Slave and Dallas Buyers Club incorporate vulgarity to authentically portray the harsh realities and psychological distress of their characters. Joker, a psychological thriller, features radicalised language reflecting the protagonist’s severe mental illness. Other films may use moderate expletives to enhance the atmosphere and aesthetic.
@@
Figure 11 shows how the probability of abusive content in Oscar-nominated films and the top 10 blockbuster films changed over the years between 1950 and 2024. The probability of abusive content is the amount of abusive content in a movie dialogue, with 10 years timespan averaged. Although the average probability of abusive content fluctuates, we observe that Oscar-nominated films are far more likely to contain abusive content than the top 10 blockbuster films from the 1950s to 1970s. Afterwards, the probability of Oscar-nominated movies is almost equal to the top 10 blockbuster films.
@@
Figure 12 presents the average abusive count (normalised count of abusive content per film) for each decade in four film categories (action, comedy, drama, thriller) between 1950 and 2024. We can observe that the abusive count in action and drama films was stable and low across all decades, indicating relatively little abusive content in both categories. We note the unusually high abusive count of thrillers in the 1950s.
@@
Figure 13 presents the frequency of abusive words between Oscar-winning movies and the top 10 blockbuster movies. Among action movies, the number of abusive words in the top 10 films is much higher than that of Oscar films, nearly 7,000. This suggests that mainstream commercial action films often contain more abusive content, possibly to cater to the audience’s need for excitement and strong emotions. The low number of abusive words in Oscar films, around 3,000, may indicate that Oscar films focus more on story depth and artistic expression rather than relying on violent elements. We observe that the two categories of comedy films had relatively similar numbers of abusive words in this category, although Oscar films are slightly higher. This may indicate that there is less violence in comedies, both commercial and Oscar movies and that they are still primarily humorous.

### v3

@@ 3.2 N-gram analysis
The N-gram model is a basic NLP method used to examine ’n’ consecutive words from a given sequence of text (Tripathy et al.,, 2016). The n-gram model has been widely used in text categorisation and provides additional information about sentiment classification (Dey et al.,, 2018) and improves the classification performance of informal text-based tasks (Kruczek et al.,, 2020). The simplicity and efficiency of the n-gram model make it the preferred choice for large vocabulary speech recognition applications (Liu et al.,, 2003). Pang et al. (Pang et al.,, 2002) employed unigram, bigram, and a combination of both unigram and bigram techniques to classify positive and negative sentiments for classification. In the framework, we used bigrams and trigrams to ascertain the frequency of word occurrences, combined with sentiment analysis.
3.3 Model Selection and Fine-tuning Strategy
We selected RoBERTa (Robustly Optimised BERT) (Liu et al.,, 2019; Zhu et al.,, 2020) as it features an enhanced training approach with larger corpora, longer text sequences, and dynamic masking patterns, which proved more effective than standard BERT when fine-tuned with the SenWave dataset.
We employed HateBERT (Caselli et al.,, 2020) for abuse detection, a BERT variant specifically designed for hate speech detection. HateBERT’s pre-training on abusive content from Reddit communities makes it particularly suitable for our movie dialogue analysis task. Our fine-tuning approach involves domain-specific training to capture unique linguistic patterns in movie dialogues across different genres and decades. This enables effective detection of contextual abuse and subtle emotional expressions typical in cinematic dialogue, significantly improving practical accuracy for our longitudinal analysis framework.
@@ 3.4 Framework
We presented a framework for abuse detection and sentiment analysis of movie subtitles that includes several sequential steps, as shown in Figure 1. Our framework uses three datasets, including the movie subtitles, the SenWave dataset for sentiment analysis and RAL-E dataset for abuse detection for refining the RoBERTa models.
In Stage 1, we downloaded and curated movie subtitles. Oscar-nominated films and the top 10 Box Office films for each year were considered the most representative films, with categories shown in Table 1.
In Stage 2, we used the NLTK (Loper and Bird,, 2002) tool to remove the special phrases, such as hashtags, emotion symbols (emojis), stop words (e.g. “the”, “an”, “you”) for each movie subtitles during data processing. Stop words were also eliminated based on specific categories to ensure accurate bigrams and trigrams for subsequent analysis.
In Stage 3, we extracted the top ten bigrams and trigrams from movie subtitles for longitudinal analysis. Movie duration changes over the decades are also reported. Furthermore, the RAL-E and SenWave datasets were processed and cleaned for fine-tuning the BERT-based model at a later stage.
In Stage 4, we obtained two RoBERTa models from HuggingFace (Wolf et al.,, 2019)  [^9] https://huggingface.co/docs/transformers/en/model_doc/roberta, fine-tuned with the SenWave and RAL-E datasets for sentiment analysis and abuse detection, respectively. The SenWave dataset enables comprehensive sentiment analysis by providing multi-label classification.
In Stage 5, we categorised the sentiments based on abuse analysis and sentiment analysis. For the abuse analysis, abusive emotions were categorised such as profanity, porn, insult, and spam. The sentiment analysis provides classifications such as denial, pessimistic, thankful, and optimistic to analyze positive and negative emotions more precisely.
In Stage 6, we analysed movie subtitles from different eras and categories using data visualisation, taking into account Oscar and Blockbuster categories along with the specific groups (action, comedy, drama, horror). Next, we compared the percentages of abusive text in each category and longitudinal sentiment polarity. Finally, we provided a subsequent comparison of a wider range of emotions, both positive and negative.
In Stage 7, we selected popular Oscar and Blockbuster movies, provided sentiment analysis and abuse detection on the entire movie timeline. This is reviewed by taking snapshots of dialogues for qualitative analysis.
@@ 3.5 Technical Details
We utilise the state-of-the-art NLP techniques with pre-trained models based on RoBERTa  [^10] https://huggingface.co/FacebookAI/roberta-base, specifically the HateBERT model (Caselli et al.,, 2020) optimised for detecting hate speech. We used the pre-trained HateBERT models from the Transformers library  [^11] https://github.com/huggingface/transformers. Given the large dataset, we implemented batch processing to efficiently make predictions. Finally, we found that RoBERTa was more efficient in refining using the SenWave dataset.
@@ 4.1 Data analysis
We use n-gram analysis to outline the changes in the top twenty words of theme and sentiment in selected films over every 20 years. In Figure 2-Panel (a), we can observe that from 1950 to 1969, positive and colloquial phrases such as ”yes sir”, ”let go” and ”good night” were frequent, and these words usually appear in contexts that express affirmation or goodwill. In addition, ”ha ha ha” and “good night good” indicate that the movies from this period are positive and affirmative emotions. Between 1970 and 1989 (Figure 2)-Panel (b), we can see that common bigrams such as ”let go” and ”yes sir” still have a dominant position and ”oh god” and ”come come” have risen significantly. These changes reflect shock, surprise and action in the film. The appearance of ”come let go” and ”go go go” shows an increase in the film’s depiction of the action. In the period 1990-2009 (Figure 2-Panel c), we can see that ”let go”, ”oh god” and ”come come” still appear more frequently in bigrams than previously. The common trigrams such as ”go go go” and ”come let go” continue to appear, but ”hey hey hey” rises in frequency. The rapid increase in word frequency indicates a significant increase in the length of subtitles in movies over the past two decades. The similarity of high-frequency words indicates that the main emotions of films during this period are the same as those of different periods in the past. During the period 2010-2024 (Figure 2-Panel d), both bigrams and trigrams contain high-frequency particles, such as ”hey hey”, ”yeah yeah”, ”whoa whoa whoa” and ”oh oh oh”. This indicates that emotions in films during this period were still predominantly positive, accompanied by more vocabulary that enhanced emotional expression.
@@ 4.2 Sentiment analysis
Model Acc. Prec. Rec. Micro F1 Ham. Loss
TF-IDF + LR 0.44 0.62 0.57 0.59 0.118
TF-IDF + SVM 0.45 0.64 0.56 0.60 0.115
BERT-base 0.48 0.68 0.62 0.65 0.102
RoBERTa-base 0.52 0.71 0.66 0.68 0.094
Table 2: Performance comparison of baseline models and RoBERTa model on SenWave dataset (English, multi-label).
We evaluated the multi-label sentiment detection on the English subset of the SenWave dataset (10,000 tweets) and divided the dataset using iterative multilabel stratification into 70% training, 10% validation, and 20% test, with a fixed random seed (42) for reproducibility. We provided the results using baseline models based on TF-IDF vector representations (Das et al.,, 2023) of the text, including unigrams and bigrams (ngram_range=(1,2)), limited to the top 50,000 features by term frequency (max_features=50,000). Logistic Regression (TF-IDF + LR) and Linear Support Vector Machine (TF-IDF + LSVM) serve as classifiers.
As shown in Table 2, the traditional baseline models perform modestly, with Subset Accuracy around 0.44–0.45 and Micro-F1 below 0.61. Fine-tuned BERT-base improves to 0.65 Micro-F1 with balanced Precision (0.68) and Recall (0.62). RoBERTa-base achieves the best overall results, reaching 0.68 Micro-F1, 0.52 Subset Accuracy, and the lowest Hamming Loss (0.094). Its higher recall (0.66) relative to BERT explains the gain, highlighting the advantage of domain-robust pre-trained models for multi-label sentiment classification.
Movies are generally highly expressive, and there is often no single emotion that can be expressed in a film. The different types of emotions captured from the subtitles depend on the type of film as well as the specific plot. Figure 4 illustrates the frequency of different emotions in movies detected by RoBERTa refined using SenWave data. We find that humour (joking) is the most frequent sentiment, which indicates that humour and lighthearted elements are essential in engaging audiences. This is followed by ”annoyed”, ”sad”, ”pessimistic”, and optimistic sentiments, which include all the selected movies from 1950 to 2024. We note that these include Oscar-nominated movies and blockbusters in four major categories. Figure 5 presents the results for the movies in the four major genres, including action, comedy, drama and thriller. We can observe that the humour sentiment is expressed the most in all categories, followed by ”annoyed”. The proposition of ”sad” is different, where drama has the most and action and thriller have the least. There is a high number of ”sad” sentiments in the ”comedy” category. Most movies would fall into two categories; for example, the movie ”Annie Hall” is an example of a movie that would fall in both comedy and drama and ”The Departed” would fall in both action and thriller categories. Therefore, the sentiments detected have overlaps across the categories.
Figure 4: Counts of different emotions for Oscar-nominated and top ten blockbuster movies from 1950-2024.
@@
We input the entire movie subtitle dataset into the RoBERTa model to generate word embeddings. A fully connected layer was then appended to the RoBERTa output, with the output dimensionality corresponding to the number of sentiment dimensions. We acquired the sentiment weights ranging from -4 to 3 using an unconstrained linear activation function, enabling a direct interpretation of each sentiment dimension’s contribution to the overall polarity, as depicted in Figure 6. This demonstrates the contribution of different emotions to the overall affective polarity score, where positive numbers indicate positive emotions and negative numbers indicate negative emotions.
Based on Figure 6, we calculated the averag sentiment weight scores for every decade as shown in Figure 7. For blockbuster movies, the weighted sentiment scores are much higher during the 1950s, 1980s, and 1990s when compared to other periods, indicating that movies had richer emotional expressions during these three decades. In the case of Oscar-nominated movies, the average weighted sentiment score has been much higher than in other periods since 2020, indicating that the Oscars are gradually increasing their focus on films with richer emotional expressions. The standard deviation values in the bar chart are substantially higher during the 1950s, 1990s, and 2000s compared to other periods, indicating notable differences in emotional expression among films in these three periods.
@@
Figure 7: Average weighted sentiment score trend for the given decades comparing Oscar and Blockbuster movies. Error bars represent the standard deviation (STD).
Different themes and styles of movies cover different emotions, and multiple emotions can coexist in the same movie. Therefore, we obtain a heatmap of the co-occurrence of emotions in Figure 3, where the darker colours show higher co-occurrence frequency. We can observe that ’Humour’ and ’Annoyed’ have a high frequency that suggests many films use humour alongside anger, e.g. situations where jokes cause discomfort. Often, satire and dark humour are employed to show societal harsh realities. Similarly, ’Joking’ and ’Sad’ co-occurred 97 times, indicating that humour is often used to alleviate sadness, a characteristic commonly observed in tragicomic films. Later decades showed more overlapping emotions such as humour mixed with annoyance, suggesting modern films use more complex emotional layering.
At the same time, some low-frequency co-occurrence emotion phrases deserve our attention. ”Thankful” and ”Humour” highlight a significant lack of immediate emotional transition or simultaneous appearance between these two states in the data. The co-occurrence frequency of ”Thankful” and ”Pessimistic” is 0, which indicates that these two emotions hardly appear and in reality, thankfulness and pessimism are almost two opposite emotions and rarely appear in the same context.
@@ 4.3 Abuse and Hate-Speech detection
Model Accuracy Precision Recall F1-score
TF-IDF + Logistic Regression 0.64 0.59 0.71 0.64
TF-IDF + SVM 0.66 0.70 0.61 0.65
BERT-base 0.72 0.75 0.69 0.72
HateBERT 0.77 0.74 0.81 0.77
Table 3: Performance comparison of baseline models and HateBERT model on RAL-E dataset.
We compared three baseline models (TF-IDF + LR, TF-IDF + LSVM, and BERT-base) and the domain-adapted HateBERT model to evaluate the effectiveness of abuse detection. The RAL-E dataset, formulated as a binary classification task (abusive vs. non-abusive), is split into 70% training, 10% validation, and 20% testing sets using stratified sampling to preserve label distribution.
The results (Table 3) show that HateBERT achieved the highest accuracy score (0.77), and TF-IDF + LR had the lowest (0.64). We focus on the F1-score as it accounts for the Precision and Recall. In terms of the F1-score, we find that HateBERT gave the best value (0.77) and TF-IDF + LR remained the weakest (0.64). Overall, these results highlight the limitations of traditional approaches and show that pre-trained language models, particularly those with domain-adaptive pretraining, provide notable gains in both overall accuracy and recall robustness. Figures 8 and 9 present the results for abusive content, demonstrating how the movie subtitles have changed over the years, influenced by social-cultural, censorship, film genres, and external factors such as epidemics.
We utilise a look-up dictionary to count the number of abusive words. Figure 8 shows the trend of the number of words and abusive words in the film subtitles for each decade. We observe that the trend of the number of total words and the number of abusive words exhibits a greater degree of similarity across different decades. After the 1970s, the proportion of abusive words to total words in film subtitles across decades has not changed significantly. Such fluctuations may be attributed to changes in social tolerance for film content, the influence of censorship, and the emergence of new film genres. It is well known that abusive language was rare before 1980 but rose significantly afterwards, reflecting censorship and greater tolerance for controversial content.
@@
Figure 9: Average movie time vs average abusive movie time per decade. Error bars represent the standard deviation (STD).
@@
Figure 10 presents trends in the amount of violent content detected in Oscar-nominated films and top 10 blockbuster films between 1950 and 2024. Abusive content has shown a gradual increase since 1980, peaking in 2014, which may reflect the growth of the film industry and the changing social and cultural acceptance of abusive content. From 1950 to 1980, the amount of abuse in Oscar-nominated and top 10 blockbuster films was relatively small. This could be due to the more conservative social and cultural requirements of film content at the time. However, abusive content gradually increased after the 1980s, and the amount of abuse in the top 10 blockbuster films almost surpassed that in Oscar-nominated movies before 2010, and has fallen significantly since then. This may indicate that Oscar-nominated films have become bolder and more diverse in subject and content, and the increase in abusive content may be an attempt to better represent complex social issues and human conflicts. Overall, we observe evolving trends in the representation and amount of abusive content in films over time, which reflect changes in social culture, the film industry, and audience preferences. Before 1980, abusive content was less and more restrained, and then increased with the diversification of film genres, especially in commercial films. Abusiveness in Oscar-nominated films has increased significantly since 2010, reflecting films’ exploration of complex themes and commercialisation.
The years with the highest and lowest number of abuse for Oscars and Blockbuster movies were selected in Figure 10, and displayed the corresponding movie titles in Table 4. Analysis of Oscar-nominated films shows a peak in abusive content in 2014 and 2020. Films like The Wolf of Wall Street and The Irishman use frequent expletives to heighten dramatic tension, emphasise the protagonists’ volatile personalities, and create a gritty atmosphere. 12 Years a Slave and Dallas Buyers Club incorporate vulgarity to authentically portray the harsh realities and psychological distress of their characters. Joker, a psychological thriller, features radicalised language reflecting the protagonist’s severe mental illness. Other films may use moderate expletives to enhance the atmosphere and aesthetic.
@@
Figure 11 shows how the probability of abusive content in Oscar-nominated films and the top 10 blockbuster films changed over the years between 1950 and 2024. The probability of abusive content is the amount of abusive content in a movie dialogue, with 10-year timespan averaged. Although the average probability of abusive content fluctuates, Oscar-nominated films are far more likely to contain abusive content than the top 10 blockbuster films from the 1950s to 1970s. Afterwards, the probability of Oscar-nominated movies is almost equal to that of the top 10 blockbuster films.
@@
Figure 12 presents the average abusive count (normalised count of abusive content per film) for each decade in four film categories (action, comedy, drama, thriller) between 1950 and 2024. The abusive count in action and drama films was stable and low across all decades, indicating relatively little abusive content in both categories. We note the unusually high abusive count of thrillers in the 1950s.
@@
Figure 13 presents the frequency of abusive words between Oscar-winning movies and the top 10 blockbuster movies. Among action movies, the number of abusive words in the top 10 films is much higher than that of Oscar-winning films, nearly 7,000. This suggests that mainstream commercial action films often contain more abusive content, possibly to cater to the audience’s need for excitement and strong emotions. The low number of abusive words in Oscar films, around 3,000, may indicate that Oscar films focus more on story depth and artistic expression rather than relying on violent elements. Two categories of comedy films had relatively similar numbers of abusive words in this category, although Oscar films are slightly higher. This may indicate that there is less violence in comedies, both commercial and Oscar movies and that they are still primarily humorous.